Air with cv = 0.723 KJ/Kgm - °R is compressed in a piston-cylinder machine according to pV1.3 = c from an initial temperature of 17°C and pressure of 1 bar, to a final pressure of 5 bars. Determine the final temperature, the heat transferred and the work.


1
Expert's answer
2021-02-22T07:31:45-0500

cv = 0.723 KJ/Kg- °R,pV1.3 = c,"T_1" =17°C and pressure of P1=1 bar, to a final pressure of ,P2=5 bars.

As per polytropic relation


"\frac{P_1}{P_2}= (\frac{T_1}{T_2})^{\frac{n}{n-1}}"

"\frac{1}{5}= (\frac{290}{T_2})^{\frac{1.3}{1.3-1}}"

"T_2=420.9 K"

work done in polytropic process

"W=\frac{nR \Delta T }{n-1}= \frac{1.3 \times 8.314\times 130.899}{.3}=4708.62 J"

Now heat transfer relation is

"\Delta Q= \Delta U+ \Delta W"

"\Delta Q= nC_vdT+W"

"\Delta Q= 0.723\times 230.9 + 4708.62"

Heat transfer is Q=4875.56 J